40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 650     Entry time: Tue Jul 8 21:58:22 2008
Author: alberto 
Type: Update 
Category: General 
Subject: Secondaty beam aligned to the IFO beam again 
Yesterday the alignment of the secondary beam to the IFO was completely lost and today I had to realign all the optics before I was able to match the two spots again. I had to reset the height of the irises and I had also to replace mirror M1 with one with a larger angular motion. Eventually I obtained the beat again. Working on the optics table I inadvertently misaligned the OSA but I didn't make in time to bring it back before the night shift people came. I'll work on that tomorrow morning.
ELOG V3.1.3-